OpenStudy (anonymous):

Frodo ran 2 2/5 miles in 7/12 of an hour. How many miles can Frodo run in one hour? A 3 17/35 B 4 4/35 C 35/144 D 20 4/7

OpenStudy (anonymous):

Will give Medal.

OpenStudy (cwrw238):

divide 2 2/5 by 7/12 hours its best to convert each fraction to the improper form first

OpenStudy (cwrw238):

2 2/5 = 10/5 + 2/5 = 12/5 no need to do anything to 7/12 to divide by 7/12 first invert it then multiply it by 12 / 5

OpenStudy (cwrw238):

weird - my posts are coming out in wrong order!!!

OpenStudy (anonymous):

ty for the help

OpenStudy (anonymous):

what the anwer